# Exploring Influence Factors on LLM Suitability for No-Code Development of End User IoT Applications

### v1

@@ 4.4. Findings
According to the experimental results, OpenAI LLMs, i.e., GPT-4o and GPT-4o-mini, and Chinese LLMs, i.e., Baidu Qianfan and Ali Qwen, demonstrate strong performance in both syntactic and semantic accuracy, making them promising candidates for LLM-based FaaS platforms. Specifically, GPT-4o surpasses the Chinese LLMs in semantic accuracy while achieving a comparable syntactic success rate. The Chinese LLMs, exhibit a clear advantage over GPT-4o-mini in both syntactic and semantic accuracy. Notably, GPT-4o-mini performs well on simple tasks but struggles with complex ones due to the lightweight factor. Other LLMs, i.e., Gemini, Copilot, DeepSeek, and Llama, face challenges in maintaining both syntactic and semantic accuracy.
@@
Translating user requirements from Chinese to English has only a slight impact on the syntactic performance of OpenAI models but affects semantic accuracy more strongly, leading to a performance drop for GPT-4o but an improvement for GPT-4o-mini. Although Gemini and Copilot have limited performance, translation improves both the syntactic and semantic accuracy of Copilot while reducing these of Gemini. The variance in prompts has little effect on Llama, which consistently fails to generate functions in both Chinese and English.
@@
Interestingly, the performance of the LLMs does not entirely align with the task complexity, i.e., they perform better on the complex task than the medium and advanced tasks, except for the setting of Chinese prompts used with GPT-4o. For results using translated English prompts, even GPT-4o fails to maintain the alignment. The reason for that is because both the medium and advanced tasks explicitly specify with three unrelated sub-tasks, each of which is relatively straightforward in terms of both user requirement description and the expected function logic. However, the semantic success rate drops significantly. While the complex task involves a more complex functional logic and uncertainty and diversity of user preferences, it performs better. Specifically, the Chinese LLMs exhibit similar performance alignment with what tasks specified, except for the medium task, where user requirements generally involve triggering more devices than the advanced task. These findings suggest that while the task complexity affects the LLM performance, existence of sub-tasks also matters. For LLMs weak on natural language understanding, the medium task is more challenging than the complex task. Additionally, a strong natural language understanding capability is essential for LLMs to generate correct results.
@@
Overall, these results underscoring the challenges that LLMs face in generating semantically correct and functional outputs.
@@ 5. Discussion
To explore the factors influencing LLM-supported no-code development platforms, we conducted experiments using LLM4FaaS, a no-code application development platform combining the power of LLM and FaaS. Our experiments examined the impact of different model choices, the language used for prompt inputs, and the role of the LLMs’ origin communities. In this section, we discuss the implications of the findings and provide insights for future research.
@@ 5.1. Model Selection Strategy
With the release of a range of diverse LLMs in recent years, new opportunities for no-code development platforms rose. However, it is challenging to maintain an overview over model performances and select the most suitable for the use case at hand. With a vast number of available LLMs, each exhibiting different strengths and weaknesses, choosing the right model requires careful consideration. In this case, we select five LLMs with different design motivations to evaluate their performance to no-code development platforms. According to the experiment results, OpenAI models, i.e., GPT-4o and GPT-4o-mini consistently outperform the other models. This is true even for the LLM designed specific for software engineering tasks, e.g., Copilot, open-source LLM optimized for efficiency and strong adaptability, e.g., Llama, and LLM with advanced reasoning capability, e.g., Gemini. We conclude the following insights on selecting model for no-code development platforms.
@@ 5.1.1. General purpose LLM suits better than coding-specific LLMs
Copilot, as a coding-specific model should excel in generating functional code snippets. However, our experimental results indicate that Copilot does not perform well in no-code development platforms. The primary reason for its poor performance is that generating a function for non-technical users, requires more than just coding expertise, it also demands strong natural language understanding in order to interpret user prompts and project context accurately. While coding-specific models excel at code comprehension and generation, they tend to favor technical inputs over natural language descriptions from non-technical users. Our experiments show that Copilot often produces pseudocode, incomplete code skeletons, or fails to generate functional outputs altogether indicating that it struggles to transfer natural language requirements into concrete functional implementations. Additionally, Copilot frequently requires further iterations with more detailed information, highlighting its limitations in natural language understanding. In contrast, general-purpose LLMs balance both natural language understanding and code generation, making them more suitable for no-code development platforms. Our findings indicate that natural language comprehension is more critical than pure coding ability in this context, allowing general-purpose models to better interpret and implement user requirements.
@@ 5.1.2. Limitations of Lightweight LLMs for Complex Tasks
While general-purpose LLMs demonstrate superior performance in no-code development, the choice of model size and capability impacts task effectiveness. For simpler tasks, even a lightweight model, i.e., GPT-4o-mini performs well, efficiently handling straightforward logic and basic automation. Its lower computational cost makes it a viable option for quick iterations and low-complexity scenarios. However, as task complexity increases, even when the generated function remains relatively short, the lightweight models struggle to maintain accuracy and coherence. Complex workflows and nuanced logic require deeper contextual understanding and more robust reasoning abilities where requires a more powerful model to consistently deliver accurate and reliable results.
@@ 5.1.3. Performance Variations Among General-Purpose LLMs
Our experiment evaluates different general-purpose LLMs, i.e., OpenAI models, Gemini, and Llama, using the same structured prompt. The results reveal significant differences in their responses and only OpenAI models prove feasible for no-code development platforms. The other models generate lengthy but invaluable outputs. Gemini tends to produce excessive additional code snippets beyond the requested function. While this broader output can be useful, for our experiment, it introduces unnecessary complexity, incurs unintended errors or warnings, and leads to higher token consumption. Rather than improving the final code generation, the additional content requires users to manually extract relevant portions, which demands technical expertise and makes it less suitable for no-code development platforms. This behavior suggests that Gemini prioritizes completeness over precision. Llama, on the other hand, places focus on analyzing project context in the prompt. Instead of directly generating the requested function, it tends to rewrite or restructure existing reference code, often deviating from the explicit intent of the prompt. This indicates that Llama prioritizes context-aware code generation, which may be beneficial in some software engineering workflows but is less aligned with the needs of no-code development.
@@ 5.1.4. Community background of LLMs matters
As our end user descriptions origin in Chinese, we explore performance of comparable Chinese LLMs from Alibaba, Baidu, and DeepSeek. The Alibaba Qwen and Baidu Qianfan models demonstrated similar performance to the OpenAI models, while DeepSeek-r1-7B performed poorly in comparison. In detail, Qianfan slightly outperformed Qwen, and both surpassed GPT-4o-mini. While GPT-4o outperformed the Chinese LLMs in terms of semantic success, the two Chinese models demonstrated comparable performance to GPT-4o in terms of semantic accuracy, emphasizing the influence of community alignment to the result accuracy. The DeepSeek model generated results similar to those of Gemini and Llama, offering a comprehensive thinking process rather than directly producing the requested function. It can be useful for developers to understand code or for educational purposes, but it is not ideal for no-code development platforms, where more targeted and functional code generation is required. The results imply that even the linguistic background of LLMs influences their performance in no-code development platforms, the design focus of the model and precise output are the key determinant of success.
@@ 5.1.5. Response time varies among LLMs
The runtime of query completion is another important factor in the choice of model. Our experiments indicated large differences between the models, however, we did not perform specific measurements on the computation time. This would be an interesting aspect for future research.
@@ 5.2. Prompt Construction Concerns
The language used in the prompt input plays a crucial role in determining the LLM’s performance in no-code development platforms. We evaluate the performance of LLM-powered no-code platform with original Chinese prompts and translated English prompts. We find that achieving optimal performance requires adapting prompt design when applying different LLMs to no-code development platforms, as each model benefits from a distinct approach to maximize performance and ensure accurate code generation. For a general-purpose model with strong natural language understanding, e.g., GPT-4o, the original user prompt tends to work best, as the model can capture nuanced details from the description. For a lightweight model, e.g., GPT-4o-mini, the optimal strategy is the opposite. For a model with advanced reasoning capabilities that generates excessive results, such as Gemini, the original user prompt is preferred as it retains more detailed information, enabling the generation of comprehensive results. Retaining the original prompt is more effective for easy tasks, while translating the prompt into English benefits complex tasks by allowing the model to focus more on code generation rather than natural language interpretation. For a coding-specific model such as Copilot, English prompts generally yield better results than original Chinese ones. When provided with Chinese input, we notice that Copilot first translates the text into English before generating code, which can dispread the focus and introduce inaccuracies. However, with English prompts, Copilot is more likely to generate pseudocode or incomplete functions, limiting its suitability for no-code development. Although different LLMs respond differently to prompt language, the overall impact on performance remains limited. Notably, GPT-4o consistently outperforms other models, making it the most reliable choice for no-code development. These findings indicate that while prompt language can influence model behavior, strong natural language understanding is the key determinant of success.

### v2

@@ 4.5. Findings
Our evaluation across eight LLMs reveals a clear divergence in their ability to satisfy both syntactic and semantic requirements for NCDP. According to the zero-shot experimental results, OpenAI models, i.e., GPT-4o and GPT-4o-mini, and some Chinese LLMs, i.e., Baidu Qianfan and Alibaba Qwen, demonstrate strong performance in both syntactic and semantic accuracy, making them promising candidates for LLM-powered NCDPs. All of them are designed as general purpose LLMs, without specialized optimization for tasks such as coding or reasoning. In contrast, LLMs designed with auxiliary capabilities, e.g., Gemini, Copilot, and DeepSeek, struggle to maintain both syntactic and semantic success. These findings suggest that NCDP requires a balanced capability among different dimensions. Specifically, GPT-4o surpasses the Chinese LLMs in semantic accuracy while the Chinese LLMs exhibit a clear advantage over GPT-4o-mini in both syntactic and semantic accuracy. This suggests that the alignment of the language of LLM training data with the language of user requirements (e.g., Chinese) can enhance performance, however, the capability of LLM itself is also crucial. Notably, a lightweight LLM, such as GPT-4o-mini, performs well on simple tasks but struggles with complex ones, indicating that the size of LLM should be considered when selecting LLMs for real-world NCDPs.
@@
The effect of prompt language (Chinese vs. translated English) reveals LLM-specific behaviors that point toward broader trends in language alignment. Translating user requirements from Chinese to English has only a slight impact on the syntactic performance of OpenAI models but affects semantic success more strongly, leading to a performance drop for GPT-4o but an improvement for GPT-4o-mini. Although Gemini and Copilot have limited performance, translation improves both the syntactic and semantic accuracy of Copilot while reducing these of Gemini. The variance in prompts has little effect on LLaMA, which consistently fails to generate functions in both Chinese and English. These mixed outcomes suggest that the impact of prompt language varies depending on the model’s design or training focus. However, overall, prompt language alone does not affect the syntactic capabilities of high-performing LLMs, to which semantic factors remain more sensitive.
@@
Interestingly, the performance of the LLMs does not entirely align with the predefined task complexity, i.e., LLMs occasionally perform better on the complex task than the medium and advanced tasks, except for the experiment setting of Chinese prompts used with GPT-4o. For results using translated English prompts, GPT-4o also fails to maintain the alignment. The possible reason for that is because both the medium and advanced tasks explicitly specify with three unrelated sub-tasks, each of which is relatively straightforward in terms of both user requirement description and the expected function logic. Despite the straightforward requirements, the semantic success rate of these tasks drops significantly. While the complex task involves a more complex functional logic and uncertainty and diversity of user preferences, it performs better. Specifically, the Chinese LLMs exhibit similar performance alignment with what tasks specified, except for the medium task, where user requirements generally involve triggering more devices than the advanced task. These findings suggest that while the task complexity affects the LLM performance, the presence of sub-tasks also plays a significant role. For LLMs weak on natural language understanding, the task containing sub-tasks is more challenging than the complex one. Additionally, a strong natural language understanding capability is essential for LLMs to generate correct results.
@@
Overall, these limitations underscore the challenges that LLMs face in generating semantically correct and functional outputs.
Impact of Few-Shot Learning
Few-shot prompting introduces measurable improvements in performance among all LLMs applied in the experiments, but its effectiveness varies across model types. For high-performing LLMs such as GPT-4o, which already achieve high syntactic success rates in zero-shot settings, while including a small amount of error cases, few-shot prompting further reduces residual errors in both syntactic and semantic perspectives. LLMs exhibiting moderate semantic performance and including more error cases, such as GPT-4o-mini and Ali-Qwen-Max, benefit more from few-shot prompting, achieving a significant improvement in syntactic and semantic success rate. LLMs with strong coding capabilities but weak zero-shot performance, i.e., Copilot, also exhibit notable gains from few-shot iterations. These LLMs tend to produce abstract pseudocode or skeletal structures in zero-shot scenarios, the few-shot feedback helps to infer functional logic and complete code structures, thus improving semantic accuracy. In contrast, for LLMs with overall weak performance and limited code generation capabilities, i.e., LLaMA and Gemini, few-shot experiments offer minimal improvements in semantic success. Additionally, some practical constraints affect few-shot experimentation. For instance, Copilot lacks a usable API interface for automated prompting, requiring manual iterations to implement few-shot settings. This limitation hinders the reproducibility and scalability of few-shot evaluation across different platforms. Baidu Qianfan has a strict input token limit, which prevents us from providing error messages as feedback in the few-shot setting.
Overall, these findings indicate that few-shot learning is effective when the LLM possesses a baseline level of task understanding and semantic alignment, but less impactful for LLMs that lack foundational capabilities.
@@ 5. Discussion
To explore the factors influencing LLM-supported NCDPs, we conducted experiments using a representative platform. Our experiments examined the impact of different model choices, the language used for prompt inputs, and the role of the LLMs’ origin communities. In this section, we discuss the implications of the findings and provide insights for future research.
@@ 5.1. Model Selection Strategy
With the release of a range of diverse LLMs in recent years, new opportunities for NCDP rose. However, it is challenging to maintain an overview over model performances and select the most suitable for the use case at hand. With a vast number of available LLMs, each exhibiting different strengths and weaknesses, choosing the right model requires careful consideration. In this case, we select five LLMs that differ in design intent, optimization strategy, scale, and open-source or proprietary to assess their suitability for NCDPs. According to the experiment results, the general-purpose LLMs with strong capability, such as OpenAI models, i.e., GPT-4o and GPT-4o-mini, consistently show strong performance among tasks in different levels of complexity and outperform the other models. This is true even for the LLM designed specific for software engineering tasks, e.g., Copilot, open-source LLM optimized for efficiency and strong adaptability, e.g., LLaMA, and LLM with advanced reasoning capability, e.g., Gemini. We conclude the following insights on selecting model for NCDPs.
@@ 5.1.1. General purpose LLM suits better than coding-specific LLMs
Copilot, as a coding-specific model, should excel in generating functional code snippets, however, our experimental results indicate that Copilot does not perform well in NCDPs. Coding-specific LLMs, e.g., Copilot, should excel in generating functional code snippets, however, our experimental results indicate that Copilot does not perform well in NCDPs with zero-shot setting. The primary reason for its poor performance is that generating a function for non-technical users requires more than just coding expertise, it also demands strong natural language understanding in order to interpret user prompts and project context accurately. While coding-specific models excel at code comprehension and generation, they tend to favor technical inputs over natural language descriptions from non-technical users. Our experiments show that Copilot, a widely used coding-specific LLM, often produces pseudocode, incomplete code skeletons, or fails to generate functional outputs altogether indicating that it struggles to transfer natural language requirements into concrete functional implementations. Additionally, Copilot frequently requires further iterations with more detailed information, highlighting its limitations in natural language understanding. Few-shot experiment results further support this observation, i.e., Copilot starts to generate actual function logic during the few-shot round, building upon the skeletal structure it produced in the zero-shot setting. In contrast, general-purpose LLMs balance both natural language understanding and code generation, making them more suitable for NCDPs. Our findings indicate that natural language comprehension is more critical than pure coding ability in the NCDP context, allowing general-purpose models to better interpret and implement user requirements.
@@ 5.1.2. Limitations of Lightweight LLMs for Complex Tasks
While general-purpose LLMs demonstrate superior performance in NCDPs, the choice of model size and capability impacts task effectiveness. For simpler tasks, even a lightweight model, i.e., GPT-4o-mini, performs well, efficiently handling straightforward logic and basic automation. Its lower computational cost makes it a viable option for quick iterations and low-complexity scenarios. However, as task complexity increases, although the generated function remains relatively short, the lightweight models struggle to maintain accuracy and coherence. Complex workflows and nuanced logic require deeper contextual understanding and more robust reasoning abilities where requires a more powerful model to consistently deliver accurate and reliable results.
@@ 5.1.3. Performance Variations Among General-Purpose LLMs
Our experiment evaluates different general-purpose LLMs, i.e., OpenAI models, Gemini, and LLaMA, using the same structured prompt. The results reveal significant differences in their responses and only OpenAI models prove feasible for NCDPs. The other models can generate lengthy but invaluable outputs, i.e., Gemini tends to produce excessive additional code snippets beyond the requested function. While this broader output can be useful, for our experiment, it introduces unnecessary complexity, incurs unintended errors or warnings, and leads to higher token consumption. Rather than improving the final code generation, the additional content requires users to manually extract relevant portions, which demands technical expertise and makes it less suitable for NCDPs. This behavior suggests that some general LLMs, e.g., Gemini prioritizes completeness over precision. Other LLMs, e.g., LLaMA, on the other hand, places focus on analyzing project context in the prompt. Instead of directly generating the requested function, these LLMs tend to rewrite or restructure existing reference code, often deviating from the explicit intent of the prompt. This indicates that these LLMs prioritize context-aware code generation, which may be beneficial in some software engineering workflows but is less aligned with the needs of NCDPs.
@@ 5.1.4. Community background of LLMs matters
As our end user descriptions origin in Chinese, we explore performance of comparable Chinese LLMs from Alibaba, Baidu, and DeepSeek. The Alibaba Qwen and Baidu Qianfan models demonstrated similar performance to the OpenAI models, while DeepSeek-r1 performed poorly in comparison. In detail, Qianfan slightly outperformed Qwen, and both surpassed GPT-4o-mini. While GPT-4o outperformed the Chinese LLMs in terms of semantic success, the two Chinese models demonstrated comparable performance to GPT-4o in terms of semantic accuracy, emphasizing the influence of community alignment to the result accuracy. The DeepSeek model generated results similar to those of Gemini and LLaMA, offering a comprehensive thinking process rather than directly producing the requested function. It can be useful for developers to understand code or for educational purposes, but it is not ideal for NCDPs, where more targeted and functional code generation is required. The results imply that although the linguistic background of LLMs influences their performance in NCDPs, the design focus of the model and precise output are the key determinant of success.
The performance of Chinese LLMs reveals that linguistic alignment can enhance LLM performance in NCDP, but cannot replace the model design focus and capacity. Since the user descriptions in our dataset are originally written in Chinese, we evaluated the behavior of several representative Chinese LLMs, each with distinct design priorities, ranging from large-scale language understanding to reasoning support and enterprise-oriented deployment. Most Chinese LLMs, e.g., Alibaba Qwen and Baidu Qianfan, demonstrate semantic accuracy comparable to top-performing English models, suggesting that aligning linguistic context with the input prompt language offers an initial advantage in understanding user intent. However, this advantage does not always translate into higher semantic success. For example, DeepSeek-R1, while designed to prioritize interpretability and step-by-step reasoning, underperforms significantly due to its limited ability to generate concrete and executable code. Instead, it tends to output general explanations or planning-oriented text, which may be valuable in educational or debugging scenarios, but fails to satisfy the direct code generation demands of NCDPs. These observations underscore that linguistic familiarity enhances model comprehension, but the primary determinant of NCDP performance lies in the design orientation and the capacity to produce precise and functional outputs.
@@ 5.1.5. Response time varies among LLMs
The runtime of query completion is another important factor in the choice of model. While we did not conduct systematic measurements of computation time, observations suggest potential usability impacts. Gemini is the fastest overall with an average response time of 9.96 seconds, but exhibits a weak functional and semantic performance. This suggests that lower latency may be associated with limited reasoning or decoding depth in underperforming models. In contrast, DeepSeek-R1 shows the longest average response time with a mean of 72.34 seconds and max of 238.39 seconds, but without proportional gains in output quality. GPT-4o and GPT-4o-mini achieved a more practical balance, with moderate average latency of 18.94 seconds and 16.82 seconds, respectively, together with high success rates. Ali-Qwen-Max has a higher latency of 28.34 seconds, which may raise deployment concerns despite the few-shot improvements. As Copilot lacks API support, making it incomparable in timing metrics. While we did not systematically evaluate response time, these incidental observations underscore its potential impact on model usability, particularly for real-time or large-scale applications. This would be an interesting aspect for future research.
@@ 5.2. Prompt Construction Concerns
The language used in the prompt input plays a crucial role in determining the LLM’s performance in NCDPs. We evaluate the performance of LLM-powered NCDP with original Chinese user requirements and translated English version. We find that achieving optimal performance requires adapting prompt design when applying different LLMs to NCDPs, as each model benefits from a distinct approach to maximize performance and ensure accurate code generation.
@@
For a general-purpose model with strong natural language understanding, e.g., GPT-4o, the original user prompt tends to work best, as the model can capture nuanced details from the description. For a lightweight model, e.g., GPT-4o-mini, the optimal strategy is the opposite. For a model with advanced reasoning capabilities that generates excessive results, e.g., Gemini, the original user prompt is preferred as it retains more detailed information, enabling the generation of comprehensive results. Retaining the original prompt is more effective for easy tasks, while translating the prompt into English benefits complex tasks by allowing the model to focus more on code generation rather than natural language interpretation. For a coding-specific model, e.g., Copilot, English prompts generally yield better results than original Chinese ones. When provided with Chinese input, we notice that Copilot first translates the text into English before generating code, which can dispread the focus and introduce inaccuracies. However, with English prompts, Copilot is more likely to generate pseudocode or incomplete functions, limiting its suitability for NCDPs. Although different LLMs respond differently to prompt language, the overall impact on performance remains limited. Notably, GPT-4o consistently outperforms other models, making it the most reliable choice for NCDPs. These findings indicate that although prompt language can influence model behavior, strong natural language understanding is the key determinant of success.
5.3. Differential Effectiveness of Few-Shot Learning across LLMs
Our experiments reveal that the few-shot prompting is effective for all applied LLMs, but the reason and adoption decision varies across LLMs. The outperforming LLMs, such as GPT-4o, shows benefit in both syntactic and semantic factors by iterating the syntactic errors with LLMs. While there are limited cases showing syntactic errors, their strong and balanced performance can resolve the remaining runtime errors and provide executable and semantically correct function code. Medium-performing LLMs, e.g., GPT-4o-mini and Ali-Qwen-Max, exhibit more pronounced improvements. These LLMs have sufficient understanding and coding capacity to recognize and fix earlier mistakes when provided with feedback, and the greater room for improvement leads to more substantial gains. The effectiveness here implies that such models are under leveraged in zero-shot setting, and still retain significant adaptability through prompting. While these models showcase substantial improvements with few-shot prompting, their semantic success results in few-shot setting mostly approaches but still falls short of the zero-shot performance of the top-performing LLMs. This suggests that their adaptability can compensate for initial weaknesses but may not fully bridge the performance gap with stronger models. The case of Copilot highlights another dimension of this dynamic. Unlike general-purpose LLMs, Copilot is specialized for code generation and likely trained on narrower but denser programming corpora. Its tendency to generate abstract pseudocode in zero-shot settings, and its reliance on few-shot settings to complete functional logic, suggests that it prioritizes template over semantic generalization. Feedback-based few-shot experiments effectively guide this refinement process, helping such models bridge the gap between high-level structure and executable logic. In contrast, LLMs with overall weak performance in the zero-shot setting, such as LLaMA and Gemini, demonstrate minimal improvement in few-shot experiments. While some syntactic errors can be addressed, their limited ability to generalize user intent prevents them from making meaningful semantic progress, indicating their unsuitability for NCDPs.
These results highlight that the practical adoption of few-shot prompting in NCDPs requires careful consideration. Although few-shot prompting can enhance syntactic and semantic success when runtime error based iteration is feasible, it comes with additional costs in terms of token consumption, inference time, and potential financial expense. For models without API support, e.g., Copilot, few-shot setups may also require manual and additional engineering effort, making them less practical in large-scale deployment. Therefore, the adoption of few-shot prompting should consider three factors: (i) the LLMs inherent capabilities in both syntax and semantics, (ii) the complexity and tolerance for errors of the target NCDP tasks, and (iii) the available computational or operational resources. In settings where high semantic accuracy is critical and feedback-based iteration is feasible, few-shot prompting offers clear advantages. However, in resource-constrained environments or for models with limited generalization ability, the trade-offs may outweigh the benefits.
5.4. Threats to Validity
To contextualize the implications of our findings on the suitability of LLMs in NCDPs, we reflect on several factors that may affect the validity of our study. In particular, we discuss potential limitations related to the applied dataset (Section 5.4.1), the experimental platform (Section 5.4.2), the selected use case (Section 5.4.3), and the long-term relevance of our observations amid ongoing LLM advancements (Section 5.4.4).
5.4.1. Applied Dataset
We use an existing dataset containing answers from 26 real users, each providing descriptions for four smart home automation tasks of varying complexity, resulting in 104 task instances in total. While a larger dataset could improve statistical generalizability, our study is exploratory and qualitative in nature, aiming to derive practically generalizable insights into the factors affecting LLM suitability in NCDP. Moreover, the applied dataset is carefully designed to reflect how real non-technical users instruct LLM according to tasks of varying complexity, which aligns well with our research focus. Therefore, we believe that this dataset is both sufficiently large to support our conclusions and well-suited to the objectives of our study.
5.4.2. Base Platform Choice
To ensure a controlled environment and isolate LLM-specific effects, we adopt LLM4FaaS as the base platform for evaluation. Notably, our study focuses on understanding the factors that influence LLM suitability in LLM-powered NCDPs, which the derived insights can be more broadly applicable than comparing specific platforms. We select LLM4FaaS because of its clean architecture, i.e., functional logic generation is the responsibility of the LLM, while infrastructure abstraction is achieved through FaaS. This design avoids additional operations or steps to achieve a reliable performance, ensuring that any performance changes resulting from modifications to LLM-related parameters can be attributed to the LLM itself. Therefore, as an end-user-oriented, LLM-powered NCDP, LLM4FaaS aligns with the goals of our study by providing a controlled and representative environment to isolate LLM-specific behaviors and derive generalizable insights. For those interested in platform-level comparisons and design details, we direct them to the LLM4FaaS paper (Wang et al., 2025).
5.4.3. Use Case Choice
Our findings are based on the smart home use case, which we believe that this domain fits well for deriving broadly applicable insights into LLM performance. It presents realistic, diverse, and often ambiguous user requirements which is an ideal testbed for evaluating how LLMs interpret and execute natural language inputs in real-world scenarios. Comparing to domains that offer focused scenarios, e.g., chatbots or simple workflows, smart home automation encompasses a wider variety of intents, including scheduling, exception handling, and device integration, reflecting the kinds of requirements often seen in end-user development contexts. While domain-specific context exists, the core challenge of translating user intent into functional logic is broadly applicable. Thus, smart home automation provides a representative and meaningful use case for evaluating the factors that influence LLM suitability in NCDPs.
5.4.4. Long-Term Relevance of LLM Impacts
While the capabilities of LLMs are expected to improve continuously, potentially mitigating some of the limitations identified in our evaluation, we argue that the underlying challenges related to design intent, community norms, and language-specific characteristics will likely persist. Also, as LLM performance improves, user expectations are expected to rise, which may shift—rather than eliminate—the boundaries of current limitations. Therefore, our insights regarding model selection remain relevant for guiding future NCDPs, particularly in contexts where human intent and domain-specific factors play a central role.